OpenStudy (anonymous):

I need Help with Pre-Cal can anyone help me? Solve using the substitution method

OpenStudy (jhannybean):

whats the question?

OpenStudy (anonymous):

1. 4x-y=9 2x+3y=-27 And there is a bracket on the whole left side of the question

OpenStudy (jhannybean):

Mind drawing it out using the "draw tool? :\

OpenStudy (jdoe0001):

4x-y=9 [2x+3y]=-27 ?

OpenStudy (jhannybean):

oh okay.

OpenStudy (anonymous):

I can try idk how to do the brackets tho but let me try

OpenStudy (anonymous):



OpenStudy (jhannybean):

oh...

OpenStudy (anonymous):

That line drawn is a bracket im not good at drawing things like this

OpenStudy (jhannybean):

in your post before you wrote 3y and now you're writing 33..... which one is it? :|

OpenStudy (anonymous):

sorry its 3y

OpenStudy (dan815):